40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 5 18:57:10 2011, Jamie, Update, SUS, More normalization of all sus controllers 
    Reply  Tue Jul 5 21:26:51 2011, rana, Update, SUS, More normalization of all sus controllers Screenshot-1.png
       Reply  Wed Jul 6 02:12:36 2011, kiwamu, Update, SUS, Re : More normalization of all sus controllers 
          Reply  Wed Jul 6 10:35:35 2011, Jamie, Update, SUS, Re : More normalization of all sus controllers 
             Reply  Wed Jul 6 15:32:32 2011, Jamie, Update, SUS, Re : More normalization of all sus controllers 
       Reply  Wed Jul 6 11:45:20 2011, Jamie, Update, SUS, More normalization of all sus controllers 
Message ID: 4942     Entry time: Tue Jul 5 21:26:51 2011     In reply to: 4941     Reply to this: 4943   4945
Author: rana 
Type: Update 
Category: SUS 
Subject: More normalization of all sus controllers 

This is getting closer, but with the whitening left OFF and the cts2um filter also OFF, none of the suspensions are working correctly. I'm shutting down all the watchdogs until someone gets around to setting the damping gains and filters correctly.

I'm attaching a screenshot of some of the problems I see so far with MC3.

I'm going to try to get the MC suspensions working OK for tonight so that we can use them for the PRMI locking work.

Update #1: None of the MC SUS DAQ channels are found by dataviewer....SUS debugging speed reduced by 10x.  Tue Jul 05 21:38:17 2011

Update #2: POS/PIT/YAW BIAS sliders now seem to work, but are ~1000x too weak to do anything.   Tue Jul 05 21:41:38 2011

 

Attachment 1: Screenshot-1.png  432 kB  Uploaded Tue Jul 5 22:34:00 2011  | Hide | Hide all
Screenshot-1.png
ELOG V3.1.3-